If I invest $6000 at 6% per year, how much mnoey should I invest at 8% per year so that the annual income from the two investments is 7%?
Answer by richwmiller(17219)

In the future submit one problem of a type.
0.08*x+0.06*6000=0.07(6000+x)
0.08*x+360=420+0.07x
0.08x-0.07x=420-360
0.01x=60
x=6000 at 8% to make 7% on 12000
check
0.08*6000+0.06*6000=0.07(6000+6000)
480+360=0.07(12000)
840=840
ok